OpenStudy (anonymous):

Please help!! Fan + Medal **Proportions** State if each pair of ratios forms a proportion 1. 4/2 and 20/6 2. 3/2 and 18/8 3. 4/3 and 16/12

OpenStudy (photon336):

Well 4/3 = 16/12 4*4 = 16 3*4 = 12 So yes that's a proportion notice how both the numerator and denominator are multiplied by the same number